OpenStudy (anonymous):

The half-life of a certain radioactive material is 32 days. An initial amount of the material has a mass of 361 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 5 days. Round your answer to the nearest thousandth.

OpenStudy (xapproachesinfinity):

the decay function is always \(f(t)=A(\frac{1}{2})^{\frac{t}{k}}\)

OpenStudy (xapproachesinfinity):

A is the initial amount

OpenStudy (anonymous):

I think its \[y=361 \left(\begin{matrix}1 \\ 2\end{matrix}\right)^{\frac{ 1 }{ 32x }}\]

OpenStudy (anonymous):

323.945 kg

OpenStudy (xapproachesinfinity):

why x on bottom?

OpenStudy (anonymous):

I couldnt figure out how to get it to the top right lol

OpenStudy (anonymous):

well the middle of the 1/32

OpenStudy (anonymous):



OpenStudy (xapproachesinfinity):

ok so your function is \(y=361(1/2)^{x/32}\)

OpenStudy (xapproachesinfinity):

ok good then

OpenStudy (anonymous):

Am i right?

OpenStudy (xapproachesinfinity):

yeah !

OpenStudy (anonymous):

Thank you for double checking for me :)

OpenStudy (xapproachesinfinity):

np!